OpenStudy (anonymous):

3+square root of (x+3)=x

OpenStudy (anonymous):

\[3+\sqrt{x+3}=x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

by squaring both sides we get: 9+x+3=\[x ^{2}\]

OpenStudy (anonymous):

\[x ^{2}-x-12=0\] \[here \sum =-1,product equals=-12\]

OpenStudy (anonymous):

so \[x ^{2}-4x+3x-12=0\] x(x-4)+3(x-4)=0 therefore x-4=0 and x+3=0 so x=4 or x=-3

OpenStudy (anonymous):

but answer shows to be {6}

OpenStudy (anonymous):

isolate that radical firs before you square both sides... subtract 3 from both sides first.

OpenStudy (anonymous):

\[\huge 3+\sqrt{x+3}=x\] \[\huge \sqrt{x+3}=x-3\] \[\huge x+3=x^2-6x+9\] \[\huge 0=x^2-7x+6\] solve from here...